# Image-based Visibility Analysis Replacing Line-of-Sight Simulation: An Urban Landmark Perspective

### v1

@@ Abstract
Visibility analysis is one of the fundamental analytics methods in urban planning and landscape research, traditionally conducted through computational simulations based on the Line-of-Sight (LoS) principle. However, when assessing the visibility of named urban objects such as landmarks, geometric intersection alone fails to capture the contextual and perceptual dimensions of visibility as experienced in the real world. The study challenges the traditional LoS-based approaches by introducing a new, image-based visibility analysis method. Specifically, a Vision Language Model (VLM) is applied to detect the target object within a direction-zoomed Street View Image (SVI). Successful detection represents the object’s visibility at the corresponding SVI location. Further, a heterogeneous visibility graph is constructed to address the complex interaction between observers and target objects. In the first case study, the method proves its reliability in detecting the visibility of six tall landmark constructions in global cities, with an overall accuracy of 87%. Furthermore, it reveals broader contextual differences when the landmarks are perceived and experienced. In the second case, the proposed visibility graph uncovers the form and strength of connections for multiple landmarks along the River Thames in London, as well as the places where these connections occur. Notably, bridges on the River Thames account for approximately 30% of total connections. Our method complements and enhances traditional LoS-based visibility analysis, and showcases the possibility of revealing the prevalent connection of any visual objects in the urban environment. It opens up new research perspectives for urban planning, heritage conservation, and computational social science.
@@ 1 Introduction
Visibility analysis is a fundamental analytical method in urban and environmental research, and is primarily conducted based on the Line of Sight (LoS) principle (Fisher-Gewirtzman et al., 2013; Cilliers et al., 2023). LoS establishes a hypothetical direct line between an observer and the target object, with the presence of an uninterrupted line indicating visibility. Owing to its simplicity and intuitive nature, LoS has been widely adopted and underpins more advanced visibility analyses, such as viewshed analysis (Tandy, 1967; Amidon and Elsner, 1968), isovist analysis (Benedikt, 1979; Batty, 2001), to map the extent and boundaries of visible space, and visibility graph analysis (Turner et al., 2001) to investigate the spatial relationship defined by visual links. Typical use cases of visibility analysis include assessing the visual impact of new constructions (Klouček et al., 2015; Alphan, 2021; Cilliers et al., 2023), evaluating the aesthetic value and visual quality of natural landscapes (Inglis et al., 2022; Swietek and Zumwald, 2023), investigating human perception and experience in public space (Luo et al., 2025; Natapov et al., 2024; Krukar et al., 2021).
@@
Despite the widespread use, visibility analysis faces two key limitations when applied in the 3D urban environment: First, its reliability relies on the completeness and resolution of the underlying spatial data models, such as building footprints with associated height attributes, 3D city models, Digital Surface Models (DSM), and Digital Elevation Models (DEM) (Wróżyński et al., 2024; Cilliers et al., 2023; Morello and Ratti, 2009; Lei et al., 2023). High‑quality and open 3D datasets are still scarce in many cities, and studies continue to rely on 10–30 m DEMs, a resolution that can misrepresent LoS and introduce substantial error (Inglis et al., 2022). At the same time, there are other urban data sources such as Street View Imagery (SVI) that are almost ubiquitous, but have not been used for such use case. Thus, it is worthwhile to investigate whether there are other urban data sources that could be used in lieu of 3D data, allowing visibility analyses to be conducted where there is no 3D data available.
@@
Second, traditional LoS‑based methods can be blind to visual context and inter‑object relationships, aspects central to human perception (Gillings and Wheatley, 2001; Déderix, 2019; Inglis et al., 2022). By treating visibility as a mere geometric intersection, they overlook the environmental context, such as lighting, vegetation, advertising boards, architectural embellishments, and the way multiple objects share a view. The challenge becomes evident when analysing named objects such as urban landmarks. Tracing an unobstructed LoS to the apex of a landmark may confirm it is “visible,” but this says little about how the landmark is framed by neighbouring buildings, how much of its recognisable silhouette emerges beyond tree canopies, or how it co‑occurs with other icons along a skyline or view corridor. In which ways landmarks are perceived in urban environment are especially emphasised in spatial planning and heritage conservation fields (Czyńska and Rubinowicz, 2019; Ashrafi et al., 2021; Talamini et al., 2023). However, without incorporating the contextual cues and inter‑object relationships, LoS analysis cannot capture the perceptual reality that planners, designers, and everyday observers actually experience.
Previous efforts to address the limitations of visibility analyses remained mainly within the LoS paradigm, focusing on enhancing the underlying data models’ geometric resolution and semantic richness. Specifically, comprehensive 3D environments have been constructed using semantically segmented point clouds and 3D modelling tools (Luo et al., 2025; Wróżyński et al., 2024; Moon et al., 2023). However, scaling such approaches to the urban level is challenging due to substantial data preparation and computational resource demands. Instead, in this study, we propose geo-located SVI as an effective and unified medium that simultaneously captures both visibility and the rich semantic information necessary for understanding real-world visual perception. We challenge the classical reliance of visibility analysis on geometry-based LoS detection.
@@
Unlike LoS analysis, which simulates the visual process based on data models, images directly capture the visibility results of potential observers in the real world. In previous practices, images often serve as supplementary reference to LoS-based visibility analysis, offering intuitive insights about both visibility and visual contexts (Greater London Authority, 2012; Talamini et al., 2023). While their potential for large-scale, high-quality visibility assessments of urban objects remains underexplored. Two recent advancements offer opportunities for quantitatively detecting LoS from images. On the one hand, the emergence of Vision-Language Models (VLM), such as CLIP (Radford et al., 2021), OWL-ViT (Minderer et al., 2022), Grounding DINO (Liu et al., 2024) have significantly lowered the barrier to extracting semantic information and open-domain objects from images. These advancements enable identifying named urban objects with distinct visual characteristics, such as unique height, architectural style, or shape. Meanwhile, the potential of SVI as an important geo-located data source is receiving more attention in urban and geographical studies (Ito et al., 2024). SVI is typically used to observe and map the foreground space of a city for less visual obstruction (Fan et al., 2025), while neglecting distant objects in the background. However, as shown in Figure 1, certain iconic urban landmarks (e.g., The Shard in London) can be occasionally observed from SVI at street level, from varying distances and urban settings. The observation reveals both the visibility of the object, commonly in the image’s background, and relevant visual contexts, e.g., in which environment the observation happens and which semantic elements or objects are shown together with the target. Furthermore, SVI’s high availability on the street network makes it a ubiquitous and interconnected visual agent, enabling it to reveal the inherent visual relationship among multiple urban objects from a systematic perspective.
@@
With the gap and opportunities, in this study, we explore the possibility for a novel image-based visibility detection method, leveraging computer vision and geo-located SVI to replace traditional LoS-based visibility analysis. The questions to answer in this study include:
@@
    Can images be quantitatively applied to replace traditional LoS-based analysis, to investigate the visibility of urban objects at a large scale?
@@ 2.2 Common Methods and Limitations for Investigating Visibility
To overcome these constraints, 3D visibility analysis evolved to simulate real-world scenes more accurately. Several studies extended isovist analysis to 3D spatial environments (Hagberg et al., 2008; Kim et al., 2019; Krukar et al., 2021). In parallel, 3D viewshed analysis has become common practice in large-scale applications such as evaluating landscape visual quality (Czyńska and Rubinowicz, 2019; Swietek and Zumwald, 2023) or assessing the visual impact of new developments (Inglis et al., 2022; Klouček et al., 2015; Alphan, 2021; Cilliers et al., 2023). 2.5D data models, such as the Digital Surface Models (DSM), Digital Elevation Models (DEM), and vector-based building data, are applied to simulate the real-world environment surfaces (Wróżyński et al., 2024). Viewshed, cumulative viewshed, and fuzzy viewshed analyses (Cilliers et al., 2023) are now standard tools in GIS platforms (e.g., ArcGIS (ESRI, 2021), QGIS (Cuckovic, 2016), GRASS (Neteler et al., 2012)). Despite their enhanced realism, 2.5D models frequently lack semantic differentiation and are restricted to representing complex 3D morphology (Pyka et al., 2022). To address this, recent studies apply true 3D data, such as LiDAR point clouds and voxel grids, to enhance detail and accuracy (Wu et al., 2021; Zhao et al., 2020; Wróżyński et al., 2024; Fujiwara et al., 2025). For instance, Czyńska and Rubinowicz (2019) employed a high‐resolution (0.5 m grid) and semantically enriched DSM to quantify the exposure levels and spatial extent of five prominent landmarks in Lublin, Poland. However, large‐scale implementation of such comprehensive data models remains computationally intensive and is often constrained by data availability.
@@ 2.3.2 Imaging Lines of Sight via Visual-Language Models
A more prevalent application of VLM is in zero-shot object detection, which allows searching and localising target objects in image space without additional pre-training or fine-tuning of the model. Well-known zero-shot object detection models include Grounding DINO (Liu et al., 2024) and OWL-ViT (Minderer et al., 2022). Both of them accept textual queries for searching the target, while OWL also allows image queries. It is assumed that urban landmarks, as distinct visual labels, can be detected naturally from SVI with a zero-shot detection model. The detection process can be regarded as ‘observation’ from machine vision at a specific SVI location. In this way, LoS towards landmarks and the meaningful visual context and connections can be identified, which defines the image-based visibility analysis method. The details of the method will be elaborated in the following sections.
@@ 3.1.1 Step 1: Calculate Relative Position and Angle
Using the planar axes convention (east $x$, north $y$), we can calculate the horizontal azimuth of landmark compared to the observer from north as:
@@ 4.1.1 SVI-based Visibility vs 3D Simulated Visibility
In the first step, we compare the image-based visibility results with the 3D simulated visibility results at SVI locations, by evaluating the classification accuracy metrics for ‘visible’ / ‘invisible’ with the same validation dataset created. We carry out 3D visibility simulation with the Python package of ‘VoxCity’ [^1] https://github.com/kunifujiwara/VoxCity (Fujiwara et al., 2025). ‘VoxCity’ is a one-stop toolset for collecting and converting open 3D urban data to voxels, and conducting 3D spatial analysis based on them. 3D visibility simulation is carried out within a 3000-meter radius from each landmark, using a uniform 3D voxel grid size of 5 × 5 × 5 m. The data models applied to generate voxels incorporate buildings, tree canopy and terrain, whose sources are detailed in Table 5 in Appendix C. We employed different data sources for each city, depending on the sources’ specific geographic coverage areas. Furthermore, to explore the underlying differences between the two visibility detection methods, we examine the proportions of key street view elements, such as buildings, greenery, and sky, at locations classified as ‘only SVI visible’, ‘only 3D visible’, ‘both visible’, and ‘both invisible’. These street view elements are derived from semantic segmentation on SVI, supported by the Python package ‘ZenSVI’ [^2] https://github.com/koito19960406/ZenSVI, developed by Ito et al. (2025).
@@ 4.2.1 Different Roles of Landmarks in Riverside Landscape
Urban design regulations often emphasise the preservation of direct visibility between landmarks and specific street locations in urban development. However, beyond the direct visibility, the long-term mode in which the landmarks are observed, and the visual relationships among multiple landmarks also hold significant preservation value. For example, a common experience applicable to the selected landmarks along the Thames River is that they are both part of the landscape and prominent viewing locations. However, little knowledge has been gained about whether a landmark serves more as a viewing point or as an observed object in urban settings. The inter-visibility relationship between landmark nodes is measured to answer the question. The differences in the roles of modern and historical landmarks are further discussed.
In urban sightseeing, beyond direct inter-visibility, people often experience multiple landmarks appearing together as part of the skyline. Some landmarks are frequently seen together, which reinforces their visual connection in the public’s perception, while others tend to appear in isolation, emphasising the unique identity of specific local spaces. To analyse this, we summarise the frequency and form in which landmarks are included in the visual co-existence relationship.
@@ 7 Conclusion
With SVI and zero-shot object detection as a new basis for identifying the classical ‘line of sight’ relationship, the study proposes a novel visibility analysis method and allows space for analysing and revealing broader visual and spatial contexts of observation. Further, a heterogeneous visibility graph is constructed to address the complex relationships and interactions among visual objects and quantify their connection strength. Using two case studies to investigate the visibility and visual relationship of urban landmarks, the study presents our methods’ reliability and wide application in urban design and heritage conservation. Specifically, the image-based visibility analysis fills the gaps in traditional 3D-based visibility methods, such as the limited availability of high-resolution 3D data, a lack of human-centred perspective, and weak reliability and interpretability. Furthermore, the proposed visibility graph showcases the possibility of revealing the prevalent connection of any visual objects in the urban environment, benefiting computational social science and urban system research.
Acknowledgments
We express our gratitude to the members of the NUS Urban Analytics Lab for their valuable discussions and insights. The first author is supported by the National University of Singapore under the President’s Graduate Fellowship. This research has been supported by Takenaka Corporation. This research is part of the project Large-scale 3D Geospatial Data for Urban Analytics, which is supported by the National University of Singapore under the Start Up Grant R-295-000-171-133. This research is part of the project Multi-scale Digital Twins for the Urban Environment: From Heartbeats to Cities, which is supported by the Singapore Ministry of Education Academic Research Fund Tier 1.
Declaration of generative AI and AI-assisted technologies in the writing process
During the preparation of this work the authors used ChatGPT in order to proofread the text. After using this tool, the authors reviewed and edited the content as needed and take full responsibility for the content of the publication.

### v2

# From Street View to Visibility Network: Mapping Urban Visual Relationships with Vision–Language Models

@@ Abstract
Visibility analysis in urban planning has traditionally relied on line-of-sight (LoS) simulations, which capture geometric occlusion. But these require accurate 3D data that is seldom available and often fail to represent how specific urban objects (e.g., landmarks) are visually encountered in real streetscapes. We reformulate object visibility assessment as an urban visual search problem in image space and take advantage of the omnipresent availability of street view imagery (SVI): given a reference image of the target object, we search for where it is visually detectable in SVI, bypassing the need for 3D or high-resolution elevation models. Specifically, a Vision Language Model (VLM) is applied to detect the target object within direction and zoom-controlled SVI. Successful detection represents the object’s visibility at the corresponding viewpoint. Further, taking traditional visibility analysis forward, a heterogeneous visibility graph is constructed to model the interaction between viewpoints and target objects. This graph enables mapping where visual connections occur, how strong they are, and how multiple landmarks become jointly connected through shared viewpoints. In a case study across six tall landmark constructions in global cities, the image-based method achieves an overall accuracy of 87% and reveals contextual differences in how landmarks are perceived and experienced in surrounding urban fabrics. In another case along the River Thames in London, the visibility graph uncovers the form and strength of multi-landmark connections and pinpoints the locations that mediate them. Bridges account for approximately 30% of all connections. Our method can serve as an effective alternative to traditional LoS-based visibility analysis, and showcases the possibility of revealing the prevalent connection of a wide range of visual objects in the urban environment. It opens up new research perspectives for urban planning, heritage conservation, and computational social science.
@@ 1 Introduction
Visibility analysis is a fundamental analytical method in urban and environmental research, and is primarily conducted based on the Line of Sight (LoS) principle (Fisher-Gewirtzman et al., 2013; Cilliers et al., 2023). LoS establishes a hypothetical direct line between an observer and the target object, with the presence of an uninterrupted line indicating visibility. Owing to its simplicity and intuitive nature, LoS has been widely adopted and underpins more advanced visibility analyses, such as viewshed analysis (Tandy, 1967; Amidon and Elsner, 1968), isovist analysis (Benedikt, 1979; Batty, 2001), to map the extent and boundaries of visible space, and visibility graph analysis to investigate the spatial relationship defined by visual links (Turner et al., 2001).
@@
Despite the widespread use, visibility analysis based on LoS faces two key limitations when applied in the 3D urban environment: First, its reliability relies on the completeness and resolution of the underlying spatial data models, such as building footprints with associated height attributes, 3D city models, Digital Surface Models (DSM), and Digital Elevation Models (DEM) (Wróżyński et al., 2024; Cilliers et al., 2023; Morello and Ratti, 2009; Lei et al., 2023). High‑quality and open 3D datasets are still scarce in many cities, and studies continue to rely on 10–30 m DEMs, a resolution that can misrepresent LoS and introduce substantial error (Inglis et al., 2022). At the same time, there are other urban data sources such as Street View Imagery (SVI) that are almost ubiquitous, but have not been used for such use case. Thus, it is worthwhile to investigate whether there are other urban data sources that could be used in lieu of 3D data, allowing visibility analyses to be conducted where there is no 3D data available.
@@
Second, traditional LoS‑based methods can be blind to visual context and inter‑object relationships, aspects central to human perception (Gillings and Wheatley, 2001; Déderix, 2019; Inglis et al., 2022). By treating visibility as a mere geometric intersection, they overlook the environmental context, such as lighting, vegetation, advertising boards, architectural embellishments, and the way multiple objects share a view. The challenge becomes evident when analysing named objects such as urban landmarks. Tracing an unobstructed LoS to the apex of a landmark may confirm it is “visible,” but this says little about how the landmark is framed by neighbouring buildings, how much of its recognisable silhouette emerges beyond tree canopies, or how it co‑occurs with other icons along a skyline or view corridor. In which ways landmarks are perceived in urban environment are especially emphasised in spatial planning and heritage conservation fields (Lopes et al., 2019; Czyńska and Rubinowicz, 2019; Ashrafi et al., 2021; Talamini et al., 2023). However, without incorporating the contextual cues and inter‑object relationships, LoS analysis cannot capture the perceptual reality that planners, designers, and everyday observers actually experience.
@@
Unlike LoS analysis, which simulates the visual process based on data models, images record the outcome of visibility in place, capturing real-world visual relationships directly. In previous practices, images often serve as supplementary reference to LoS-based visibility analysis, offering intuitive insights about both visibility and visual contexts (Greater London Authority, 2012; Talamini et al., 2023). While their potential for large-scale, high-quality visibility assessments of urban objects remains underexplored. Two recent advancements offer opportunities for quantitatively detecting LoS from images. First, the emergence of Vision-Language Models (VLM), such as CLIP (Radford et al., 2021), OWL-ViT (Minderer et al., 2022, 2024), Grounding DINO (Liu et al., 2024) have significantly lowered the barrier to extracting semantic information and open-domain objects from images. These advancements enable identifying named urban objects with distinct visual characteristics, such as unique height, architectural style, or shape. Second, the potential of SVI as an important geo-located data source is receiving more attention in urban and geographical studies (Ito et al., 2024). SVI is typically used to observe and map the foreground space of a city for less visual obstruction (Fan et al., 2025), while neglecting distant objects in the background. However, as shown in Figure 1, certain iconic urban landmarks (e.g., The Shard in London) can be occasionally observed from SVI at street level, from varying distances and urban settings. Furthermore, because SVI is widely available across urban streets, it forms a dense and interconnected set of geo-referenced viewpoints, allowing visibility analysis to move beyond isolated locations toward a city-scale visibility network.
@@
Given the gaps and opportunities, this study explores an image-based visibility detection method that leverages computer vision and geo-located SVI, offering a data-driven alternative to conventional LoS-based visibility analysis. The questions to answer in this study include:
@@
    Can images be quantitatively applied as a scalable alternative to 3D models, to investigate the visibility of urban objects?
@@ 2.2 Common Methods and Limitations for Investigating Visibility
To overcome these constraints, 3D visibility analysis evolved to simulate real-world scenes more accurately. Several studies extended isovist analysis to 3D spatial environments (Hagberg et al., 2008; Kim et al., 2019; Krukar et al., 2021). In parallel, 3D viewshed analysis has become common practice in large-scale applications such as evaluating landscape visual quality (Czyńska and Rubinowicz, 2019; Swietek and Zumwald, 2023) or assessing the visual impact of new developments (Inglis et al., 2022; Klouček et al., 2015; Alphan, 2021; Cilliers et al., 2023). 2.5D data models, such as the Digital Surface Models (DSM), Digital Elevation Models (DEM), and vector-based building data, are applied to simulate the real-world environment surfaces (Wróżyński et al., 2024). Viewshed, cumulative viewshed, and fuzzy viewshed analyses (Cilliers et al., 2023) are now standard tools in GIS platforms (e.g., ArcGIS (ESRI, 2021), QGIS (Cuckovic, 2016), GRASS (Neteler et al., 2012)). Despite their enhanced realism, 2.5D models frequently lack semantic differentiation and are restricted to representing complex 3D morphology (Pyka et al., 2022). To address this, recent studies apply true 3D data, such as LiDAR point clouds and voxel grids, to enhance detail and accuracy (Kim et al., 2019; Wu et al., 2021; Zhao et al., 2020; Wróżyński et al., 2024; Fujiwara et al., 2026). For instance, Czyńska and Rubinowicz (2019) employed a high‐resolution (0.5 m grid) and semantically enriched DSM to quantify the exposure levels and spatial extent of five prominent landmarks in Lublin, Poland. However, large‐scale implementation of such comprehensive data models remains computationally intensive and is often constrained by data availability.
@@ 2.3.2 Imaging Lines of Sight via Visual-Language Models
A more prevalent application of VLM is in zero-shot object detection, which allows searching and localising target objects in image space without additional pre-training or fine-tuning of the model. Well-known zero-shot object detection models include Grounding DINO (Liu et al., 2024) and OWL-ViT (Minderer et al., 2022, 2024). Both of them accept textual queries for searching the target, while OWL also allows image queries. It is assumed that urban landmarks, as distinct visual labels, can be detected naturally from SVI with a zero-shot detection model. The detection process can be regarded as ‘observation’ from machine vision at a specific SVI location. In this way, LoS towards landmarks and the meaningful visual context and connections can be identified, which defines the image-based visibility analysis method. The details of the method will be elaborated in the following sections.
@@ 3.1.1 Step 1: Calculate Relative Position and Angle
Using the planar axes convention (east $x$, north $y$), we can calculate the horizontal azimuth of the landmark compared to the observer from north as:
@@ 4.1.1 SVI-based Visibility vs 3D Simulated Visibility
In the first step, we compare the image-based visibility results with the 3D simulated visibility results at SVI locations, by evaluating the classification accuracy metrics for ‘visible’ / ‘invisible’ with the same validation dataset created. We carry out 3D visibility simulation with the Python package of ‘VoxCity’ [^1] https://github.com/kunifujiwara/VoxCity (Fujiwara et al., 2026). ‘VoxCity’ is a one-stop toolset for collecting and converting open 3D urban data to voxels, and conducting 3D spatial analysis based on them. 3D visibility simulation is carried out within a 3000-meter radius from each landmark, using a uniform 3D voxel grid size of 5 × 5 × 5 m. The data models applied to generate voxels incorporate buildings, tree canopy and terrain, whose sources are detailed in Table 5 in Appendix C. We employed different data sources for each city, depending on the sources’ specific geographic coverage areas. Furthermore, to explore the underlying differences between the two visibility detection methods, we examine the proportions of key street view elements, such as buildings, greenery, and sky, at locations classified as ‘only SVI visible’, ‘only 3D visible’, ‘both visible’, and ‘both invisible’. These street view elements are derived from semantic segmentation on SVI, supported by the Python package ‘ZenSVI’ [^2] https://github.com/koito19960406/ZenSVI, developed by Ito et al. (2025).
@@ 4.2.1 Different Roles of Landmarks in Riverside Landscape
A common experience applicable to the selected landmarks along the Thames River is that they are both part of the landscape and prominent viewing locations. However, little knowledge has been gained about whether a landmark serves more as a viewing point or as an observed object in urban settings. The inter-visibility relationship between landmark nodes is measured to answer the question. The differences in the roles of modern and historical landmarks are further discussed.
In urban sightseeing, beyond direct inter-visibility, people often experience multiple landmarks appearing together Lopes et al. (2019). Some landmarks are frequently seen together, which reinforces their visual connection in the public’s perception, while others tend to appear in isolation, emphasising the unique identity of specific local spaces. To analyse this, we summarise the frequency and form in which landmarks are included in the visual co-existence relationship.
@@ 7 Conclusion
With SVI and zero-shot object detection as a new basis for identifying the classical Line of Sight (LoS) relationship, the study proposes a novel visibility analysis method and allows space for analysing and revealing broader visual and spatial contexts of observation. Further, a heterogeneous visibility graph is constructed to address the complex relationships and interactions among visual objects and quantify their connection strength. Using two case studies to investigate the visibility and visual relationship of urban landmarks, the study presents our methods’ reliability and wide application in urban design and heritage conservation. Specifically, the image-based visibility analysis fills the gaps in traditional 3D-based visibility methods, such as the limited availability of high-resolution 3D data, a lack of human-centred perspective, and weak reliability and interpretability. Furthermore, the proposed visibility graph showcases the possibility of revealing the prevalent connection of any visual objects in the urban environment, benefiting computational social science and urban system research.